As a result of improvements in product engineering, United Automation is able to sell one of its two milling machines. Both machines perform the same function but differ in age. The newer machine could be sold today for $63,500. Its operating costs are $21,800 a year, but in five years the machine will require a $19,100 overhaul. Thereafter operating costs will be $30,900 until the machine is finally sold in year 10 for $6,350.

The older machine could be sold today for $25,900. If it is kept, it will need an immediate $24,500 overhaul. Thereafter operating costs will be $34,000 a year until the machine is finally sold in year 5 for $6,350.   

Both machines are fully depreciated for tax purposes. The company pays tax at 35%. Cash flows have been forecasted in real terms. The real cost of capital is 13%.

a. Calculate the equivalent annual costs for selling the new machine and for selling the old machine. (Do not round intermediate calculations. Enter your answers as a positive value rounded to 2 decimal places.)

Equivalent Annual Cost
Sell new machine $
Sell old machine $
Answer #1

B C F G H I J K D E 23 4 5 6 9 10 63500 21800 21800 21800 21800 30900 30900 30900 30900 30900 21800 19100 -22225 41275 -7630

=B1+1 =D1+1 =E1+1 =F1+1 6141 =H11 -11+1 =J1+1 =K1+1 21800 21800 121800 121800 3090030900309003090030900 1 Ye: 0 2 CO 63500 3